OpenStudy (anonymous):

Find the slope of the function's graph at the given point. Then find an equation for the line tangent to the graph there. h(t)= t^3+3t at (1,4)

OpenStudy (anonymous):

calculus class right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

take the derivative of \(h(t)=t^3+3t\) by the power rule what do you get?

OpenStudy (anonymous):

\[3t ^{2}+3\]

OpenStudy (anonymous):

ok good, so \(h'(t)=3t^2+3\) to find your slope at \((1,4)\) compute \(h'(1)\)

OpenStudy (anonymous):

then use the point slope formula to find the equation of the line through \((1,4)\) with the slope you find via \(h'(1)\)

OpenStudy (anonymous):

so the slope is 6?

OpenStudy (anonymous):

yes, the derivative is a formula for finding the slope of the tangent line

OpenStudy (anonymous):

then \(y-y_1=m(x-x_1) \) with \(m=6,x_1=1,y_1=4\) gives you the equation of the line

OpenStudy (anonymous):

so y=6x+10

OpenStudy (anonymous):

that is not what i get

OpenStudy (anonymous):

\[y-4=6(x-1)\]

OpenStudy (anonymous):

you might have put \(6(x+1)\) by mistake

OpenStudy (anonymous):

oh I did, now I got it

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Thanks so much!